L-1{ s / (s2+1)2 }


1
Expert's answer
2021-09-06T19:14:36-0400

"L^{-1}[\frac{s}{(s^2+1)^2}]=\frac{d}{dt}L^{-1}[\frac{1}{(s^2+1)^2}]="

"\frac{d}{dt}[\frac{1}{2}sin(t)-\frac{1}{2}tcos(t)]=\frac{1}{2}tsin(t)."